Z is a standard normal random variable, then P =...

a. P(Z < 1.37) =

b. P(Z > −1.51) =

c. P(−1.031 < Z < 1.92) =

d. P(0.00 < Z < 1.79) =

e. (A-Grade) P(Z = 0.518) =
Answer #1

***SE From standard normal distribu normal distribution table:- (a) P(Z <1-37) = 0.9146 (b) P(z) - 1051) = 1 - P(Z <-1.51) =